# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ 3.2 Continual Training
This kind of training method utilizes existing monolingual or multilingual models for enhancing the specific languages ability (e.g., Chinese-LLaMA Cui et al. (2023a), Chinese-Alpaca Cui et al. (2023a), Chinese-Mixtral HIT-SCIR (2024), CPM-2 Zhang et al. (2021), LLaMAntino Basile et al. (2023), FinGPT Luukkonen et al. (2023), Sabiá Pires et al. (2023), Bode Garcia et al. (2024)) or extending to a wider range of languages, such as BigTrans Yang et al. (2023a), Glot500 ImaniGooghari et al. (2023) and others Ebrahimi and Kann (2021); Alabi et al. (2022); Muennighoff et al. (2022b). A branch of work is to directly continually train the foundation models without any other techniques Luukkonen et al. (2023); Yang et al. (2023a); Alabi et al. (2022). For instance, Muennighoff et al. Muennighoff et al. (2022b) introduce xP3, a corpus consisting of tasks in 46 languages, and utilizes a cross-lingual multitask fine-tuning approach to fine-tune the BLOOM and mT5 on the newly created corpora, along with the English-only P3 corpus, resulting in the production of BLOOMZ and mT0Z. Considering the importance of vocabulary for multilingual models, prior work employs vocabulary extension to improve performance for low-resource languages ImaniGooghari et al. (2023). Another branch of work is to introduce external parameters and strategies to learn new languages for multilingual models Ebrahimi and Kann (2021). For instance, Chinese-LLaMA Cui et al. (2023a) integrates an additional 20K Chinese tokens into the original vocabulary to improve the Chinese ability, then employs the Low-Rank Adaptation (LoRA) approach to facilitate efficient training on Chinese corpus. CPM-2 Zhang et al. (2021) divides the pre-training into three phases, including the monolingual pre-training, bilingual pre-training, and mixture-of-expert pre-training. Multi-phase continual training, incorporating knowledge inheritance, offers a substantial reduction in computation costs compared to training from scratch. Moreover, Wang et al. Wang et al. (2022a) systematically investigate strategies aimed at reducing dependence on traditional language resources by leveraging bilingual lexicons to fine-tune pre-trained multilingual models for underrepresented languages.
@@ 3.3 Limitations and Future Directions on Training Paradigm
Knowledge conflict. As the foundation model expands the supported languages, its multilingual knowledge accumulates into the model through continuous new data available. The newly acquired knowledge from the new data would interfere with old knowledge stored in the parameters, and thus lead to a rapid decline in model performance. This is known as catastrophic forgetting Kirkpatrick et al. (2017), which might result in even completely forgetting previously learned knowledge.
Knowledge type. The existing multilingual NLP datasets, such as XNLI Conneau et al. (2018) and XQuAD Artetxe et al. (2020), are constructed based on the translated texts from the original English version datasets Rajpurkar et al. (2016); Yin et al. (2021). This is because of the high cost of obtaining multilingual data, especially in languages with limited annotations/translators. As a result, the essence of existing datasets lies in translation-based aligned knowledge, lacking cultural and domain knowledge in non-English contexts. This type of knowledge learned by the models reduces the distribution shift ability in the real-world scenario, leading to sub-optimal performance in specialized fields (e.g., finance and legislation).
@@
  • Training Strategies. LLMs are data-driven and follow the scaling law which suggests that increasing the model size can guarantee better generalization, higher accuracy, and improved ability to capture complex patterns from data. However, such a paradigm results in the existing training strategies primarily focusing on expanding diverse data, while overlooking other multilingual training techniques, such as sampling algorithms Arivazhagan et al. (2019); Wang et al. (2020); Wu et al. (2021), training objective optimization Yang et al. (2021); Pan et al. (2021), representation space Oncevay et al. (2020); Cheng et al. (2022); Stap et al. (2023), and others Zhang et al. (2020); Wang and Zhang (2022). The potential of these traditional multilingual training strategies has not been further explored in the current data effort paradigm in existing multilingual literature.
  • Architecture. Existing LLMs mainly rely on monolingual backbone models, such as the standard Transformer, rather than undergoing architectural changes specifically targeting multilingual capabilities. For the other specific tasks, for example, LLMs designed for long texts introduce variations in positional encoding Beltagy et al. (2020); Su et al. (2024), and those designed for multi-modal tasks enhance alignment between different modalities Peng et al. (2023b); Liu et al. (2024b) have been explored in some extend. Some advanced architectures like Mamba Gu and Dao (2023) have also shown remarkable performance and may become the backbone architecture for the next generation of LLMs. Thus, for multilingual tasks, it is essential to tailor the architecture with specific variations rather than merely augmenting data on the standard Transformer.
  • Sustainability. When trying to utilize new data to enhance existing LLMs, a practical approach is to allow LLMs to continually learn from the updated data, akin to how the human brain operates, instead of creating new models Huang et al. (2022, 2023a, 2023b). Essentially, when new language data comes up, we aim to continually extend the language support of LLMs and improve their corresponding language capabilities while preserving the performance of those languages that the model has already demonstrated good performance, as forming a lifelong/incremental learning paradigm.
@@ 4.1 Direct Inference in Multilingual Models
Table 2: The performance of LLMs using different inference strategies in the multilingual scenario. The best score is highlighted in bold. $\dagger$” indicates the results are quoted from Intrator et al. (2024).
@@
Recently, significant advancements have been made in enhancing the multilingual capabilities of LLMs, greatly increasing the number of languages covered and the performance in non-English inference. Advanced models such as GPT-4 Achiam et al. (2023) and PaLM-2 Anil et al. (2023) demonstrate remarkable multilingual capabilities and support hundreds of languages. The support of multiple languages enables direct multilingual inference to become feasible and come up with a current focal point of LLM research. As shown in Table 2, we investigate the comparison between direct inference and the pre-translation method which translates the prompt into a high-resource language (e.g., English and Chinese) before inference. The direct inference achieves better performance compared to the pre-translation method based on both PaLM2-S and PaLM2-L. The result demonstrates that these two LLMs have multilingual capabilities without the need to use English as a pivot for other language tasks. Moreover, without the requirement of the translation step, the direct inference approach reduces computational overhead and simplifies the processing pipeline with higher efficiency. The observations also confirm the benefits of direct inference, including the preservation of linguistic authenticity, enhanced processing efficiency, and improved performance in low-resource languages.
@@ 4.2 Pre-Translation Inference
The direct inference may not work for all LLMs, depending on their multilingual capacities. The existing LLMs usually perform better on those high-resource languages than the low-resource ones, because of the imbalance ratio within the training data. To enhance the performance on low-resource languages, pre-translation inference standardizes the input with various languages by translating them into a pivot high-resource language (e.g., English or Chinese) before querying the LLMs, which is based on the proficiency of the pivot language within the LLMs Huang et al. (2023c); Qin et al. (2023); Etxaniz et al. (2023). However, the guarantee of this method relies on high-quality translation services available, which are not necessarily true for most languages. As a result, the translation errors would accumulate and the final output of LLMs would become wrong. Besides, translation on a pivot language obscures the cultural and linguistic nuances of the original text, which might also lead to inaccurate results. As an exploratory work, Liu et al. Liu et al. (2024a) explore the performance comparison between direct inference using native languages and inference after translating into English for multilingual tasks. Although translation can enhance the multilingual reasoning performance of English-centric LLMs, for high-resource languages and advanced LLMs, reasoning in native languages is more effective. Besides, pre-translating to English is a practical approach in terms of current LLMs as they are predominantly trained on English data from their pilot experimental results. However, with the development of direct multilingual inference, the intermediary step may not be necessary, which could allow more authentic interactions with LLMs under multilingual scenarios.
@@ 4.3 Multilingual CoT
The Chain of Thought (CoT) is an effective approach to enhance the performance of LLMs in complex reasoning Sap et al. (2020); Yu et al. (2023a); Liu et al. (2023d), which has been extensively explored in existing studies, mainly focusing on English Wang et al. (2022b); Lyu et al. (2023); Wei et al. (2022a). Liu et al. Liu et al. (2024a) propose several inference strategies to investigate the effectiveness of prompting the LLMs in the multilingual scenario. For instance, “Native-CoT” indicates that asking questions in the native language with the instruction “Let’s think step by step.” is translated into the native language. “En-CoT” indicates that asking questions in the native language but instructing reasoning in English with “Let’s think step by step in English.” and “XLT” Huang et al. (2023c) indicates that translating questions into English and solving them step-by-step. “Trans” indicates that using the translation systems to convert questions into English and then solving them step-by-step. The experimental results demonstrate that the overall performance of the CoT instruction in English (En-CoT) is better than the one with instruction in native languages (Native-CoT). The results demonstrate that the CoT is still effective while using English to form the instructions for non-English queries. On the other hand, multilingual CoT attempts to enhance the reasoning capabilities of LLMs across multiple languages Shi et al. (2022a). The multilingual CoT approach is especially beneficial for complex reasoning tasks deeply embedded in specific cultural contexts, enabling more natural and intuitive problem-solving Kim et al. (2023a); Suzgun et al. (2023). The common practice of multilingual CoT is to prompt the LLMs to establish a step-by-step reasoning process in the original language of queries, which can preserve linguistic and cultural nuances. In addition, the results corresponding to different sizes of LLaMA-2-Chat attend to examine the impact of scaling law with different multilingual inference strategies.
@@
The understanding of the syntax and semantics of various languages, such as idiomatic expressions and culture-specific references is challenging for existing LLMs. A recent study Chai et al. (2024) introduces a cross-lingual CoT reasoning instruction fine-tuning framework, which randomly replaces language fragments with those from low-resource languages, and mixes the original and target languages within a single query. Besides, this study creates multilingual CoT instruction training data, which can be used to supervise fine-tuning LLMs to reduce the performance gap across different languages. However, developing datasets for training and evaluating multilingual CoT capabilities requires a robust representation of linguistic diversity and detailed language-specific nuances, which should be further addressed.
4.4 Multilingual Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) is a methodology that integrates text generation with external knowledge retrieval, dynamically enhancing the quality of model response and accuracy by accessing relevant information Gao et al. (2023a); He et al. (2023a). This approach enables the model to utilize up-to-date or specialized knowledge in text generation, thereby increasing its practicality and reliability. A main branch of the multilingual RAG adopts to retrieve knowledge from the open domain and applies it to the in-context (i.e., augmented prompt) Zhang et al. (2023b); Shi et al. (2022b); Agrawal et al. (2023); Li et al. (2023a, b); Winata et al. (2023a); Garcia et al. (2023); Ramos et al. (2023); Kim et al. (2023b). In particular, Thakur et al. Thakur et al. (2024) propose NoMIRACL, a dataset across 18 languages, to evaluate the hallucination of LLM when given a piece of text in external retrieved knowledge. It consists of two parts, the non-relevant and the relevant subset, corresponding to the question and the passage which are non-related and related, respectively. The model is expected to answer “I don’t know” in the non-relevant subset, otherwise it is determined as a hallucination. In addition, hallucinations and off-target issues occur when incorporating LLMs with low-resource machine translation Sennrich et al. (2023), where the RAG can mitigate these issues via improving the translation quality for low-resource directions Vernikos and Popescu-Belis (2024); Fu et al. (2024); Zeng et al. (2024); He et al. (2024). Overall, the RAG can effectively facilitate LLMs to generate more reliable responses, alleviating the issues of hallucination and factual error without fine-tuning. However, achieving substantial enhancements solely through the RAG method for low-resource languages, where LLMs struggle, poses a considerable challenge. Meanwhile, it is also challenging to build retrievers that can be used in low-resource languages Conia et al. (2023).
@@ 4.5 Code-Switching
Module Model SA Aguilar et al. (2020) LID Aguilar et al. (2020) MT Srivastava and Singh (2022) Summarization Mehnaz et al. (2021)
@@
Code-switching refers to the phenomenon where communicators switch between two or more languages during linguistic interactions, based on contextual needs. This phenomenon is common in bilingual or multilingual communities, especially in spoken communication Zhang et al. (2023g); Huang et al. (2024). For example, in conversations involving Chinese-English or Spanish-English direction, speakers may freely switch between languages depending on the fluency, precision of expression, or emotional needs of the dialogue Doğruöz et al. (2021); Winata et al. (2021); Aguilar et al. (2020). Solving code-switching texts is an important and challenging task as the language IDs are not specified before LLMs inference Zhang et al. (2023c); Koto et al. (2024).
@@
As shown in Table 3, we investigate the performance of LLMs on code-switching tasks. The results demonstrate that open-source LLMs cannot deal with the code-switching problem that needs to process multiple languages simultaneously under a single query. Multiple languages increase the complexity of processing and understanding language for models and LLMs need to continue to be supervised fine-tuned to accommodate simultaneous processing in this scenario. In addition, to improve the ability to deal with code-switching texts, researchers develop new pre-training techniques to enhance the understanding capacity of this linguistic phenomenon in LLMs Peng et al. (2023a). For instance, Das et al. Das et al. (2023) modify the MLM Devlin et al. (2018) to enhance model reasoning performance in code-switching environments. By employing code-switching corpora, it masks tokens at the boundaries between two languages, thus forcing the model to learn semantic nuances at the points of language transition. Additionally, this work introduces modifications to the model structure by adding residual connections from intermediate layers to the final layer. An auxiliary loss function based on the representations of intermediate layers is also proposed, which enhances the cross-lingual ability to understand multiple languages in a single sentence. However, the cost of annotated code-switching data is expensive and it is challenging to investigate the non-parametric or semi-parametric method to adapt the code-switching problems.

### v2

@@ 3.2 Continual Training
This kind of training method utilizes existing monolingual or multilingual models for enhancing the specific languages ability (e.g., Chinese-LLaMA [299], Chinese-Alpaca [299], Chinese-Mixtral [300], CPM-2 [301], LLaMAntino [302], FinGPT [303], Sabiá [304], Bode [305]) or extending to a wider range of languages, such as BigTrans [17], Glot500 [306] and others [307, 308, 309]. A branch of work is to directly continually train the foundation models without any other techniques [303, 17, 308]. For instance, Muennighoff et al. [309] introduce xP3, a corpus consisting of tasks in 46 languages, and utilizes a cross-lingual multitask fine-tuning approach to fine-tune the BLOOM and mT5 on the newly created corpora, along with the English-only P3 corpus, resulting in the production of BLOOMZ and mT0Z. Considering the importance of vocabulary for multilingual models, prior work employs vocabulary extension to improve performance for low-resource languages [306]. Another branch of work is to introduce external parameters and strategies to learn new languages for multilingual models [307]. For instance, Chinese-LLaMA [299] integrates an additional 20K Chinese tokens into the original vocabulary to improve the Chinese ability, then employs the Low-Rank Adaptation (LoRA) approach to facilitate efficient training on Chinese corpus. CPM-2 [301] divides the pre-training into three phases, including the monolingual pre-training, bilingual pre-training, and mixture-of-expert pre-training. Multi-phase continual training, incorporating knowledge inheritance, offers a substantial reduction in computation costs compared to training from scratch. Moreover, Wang et al. [310] systematically investigate strategies aimed at reducing dependence on traditional language resources by leveraging bilingual lexicons to fine-tune pre-trained multilingual models for underrepresented languages.
@@ 3.3 Limitations and Future Directions on Training Paradigm
Knowledge conflict. As the foundation model expands the supported languages, its multilingual knowledge accumulates into the model through continuous new data available. The newly acquired knowledge from the new data would interfere with old knowledge stored in the parameters, and thus lead to a rapid decline in model performance. This is known as catastrophic forgetting [25], which might result in even completely forgetting previously learned knowledge.
Knowledge type. The existing multilingual NLP datasets, such as XNLI [256] and XQuAD [257], are constructed based on the translated texts from the original English version datasets [311, 312]. This is because of the high cost of obtaining multilingual data, especially in languages with limited annotations/translators. As a result, the essence of existing datasets lies in translation-based aligned knowledge, lacking cultural and domain knowledge in non-English contexts. This type of knowledge learned by the models reduces the distribution shift ability in the real-world scenario, leading to sub-optimal performance in specialized fields (e.g., finance and legislation).
@@
  • Training Strategies. LLMs are data-driven and follow the scaling law which suggests that increasing the model size can guarantee better generalization, higher accuracy, and improved ability to capture complex patterns from data. However, such a paradigm results in the existing training strategies primarily focusing on expanding diverse data, while overlooking other multilingual training techniques, such as sampling algorithms [313, 314, 315], training objective optimization [316, 317], representation space [318, 319, 320], and others [238, 321]. The potential of these traditional multilingual training strategies has not been further explored in the current data effort paradigm in existing multilingual literature.
  • Architecture. Existing LLMs mainly rely on monolingual backbone models, such as the standard Transformer, rather than undergoing architectural changes specifically targeting multilingual capabilities. For the other specific tasks, for example, LLMs designed for long texts introduce variations in positional encoding [322, 323], and those designed for multi-modal tasks enhance alignment between different modalities [324, 325] have been explored in some extend. Some advanced architectures like Mamba [326] have also shown remarkable performance and may become the backbone architecture for the next generation of LLMs. Thus, for multilingual tasks, it is essential to tailor the architecture with specific variations rather than merely augmenting data on the standard Transformer.
  • Sustainability. When trying to utilize new data to enhance existing LLMs, a practical approach is to allow LLMs to continually learn from the updated data, akin to how the human brain operates, instead of creating new models [297, 263, 265]. Essentially, when new language data comes up, we aim to continually extend the language support of LLMs and improve their corresponding language capabilities while preserving the performance of those languages that the model has already demonstrated good performance, as forming a lifelong/incremental learning paradigm.
@@ 4.1 Direct Inference in Multilingual Models
Table 2: The performance of LLMs using different inference strategies in the multilingual scenario. The best score is highlighted in bold. $\dagger$” indicates the results are quoted from [63].
@@
Recently, significant advancements have been made in enhancing the multilingual capabilities of LLMs, greatly increasing the number of languages covered and the performance in non-English inference. Advanced models such as GPT-4 [2] and PaLM-2 [41] demonstrate remarkable multilingual capabilities and support hundreds of languages. The support of multiple languages enables direct multilingual inference to become feasible and come up with a current focal point of LLM research. As shown in Table 2, we investigate the comparison between direct inference and the pre-translation method which translates the prompt into a high-resource language (e.g., English and Chinese) before inference. The direct inference achieves better performance compared to the pre-translation method based on both PaLM2-S and PaLM2-L. The result demonstrates that these two LLMs have multilingual capabilities without the need to use English as a pivot for other language tasks. Moreover, without the requirement of the translation step, the direct inference approach reduces computational overhead and simplifies the processing pipeline with higher efficiency. The observations also confirm the benefits of direct inference, including the preservation of linguistic authenticity, enhanced processing efficiency, and improved performance in low-resource languages.
@@ 4.2 Pre-Translation Inference
The direct inference may not work for all LLMs, depending on their multilingual capacities. The existing LLMs usually perform better on those high-resource languages than the low-resource ones, because of the imbalance ratio within the training data. To enhance the performance on low-resource languages, pre-translation inference standardizes the input with various languages by translating them into a pivot high-resource language (e.g., English or Chinese) before querying the LLMs, which is based on the proficiency of the pivot language within the LLMs [65, 327, 328]. However, the guarantee of this method relies on high-quality translation services available, which are not necessarily true for most languages. As a result, the translation errors would accumulate and the final output of LLMs would become wrong. Besides, translation on a pivot language obscures the cultural and linguistic nuances of the original text, which might also lead to inaccurate results. As an exploratory work, Liu et al. [64] explore the performance comparison between direct inference using native languages and inference after translating into English for multilingual tasks. Although translation can enhance the multilingual reasoning performance of English-centric LLMs, for high-resource languages and advanced LLMs, reasoning in native languages is more effective. Besides, pre-translating to English is a practical approach in terms of current LLMs as they are predominantly trained on English data from their pilot experimental results. However, with the development of direct multilingual inference, the intermediary step may not be necessary, which could allow more authentic interactions with LLMs under multilingual scenarios.
@@ 4.3 Multilingual CoT
The Chain of Thought (CoT) is an effective approach to enhance the performance of LLMs in complex reasoning [329, 330, 331], which has been extensively explored in existing studies, mainly focusing on English [332, 333, 334]. Liu et al. [64] propose several inference strategies to investigate the effectiveness of prompting the LLMs in the multilingual scenario. For instance, “Native-CoT” indicates that asking questions in the native language with the instruction “Let’s think step by step.” is translated into the native language. “En-CoT” indicates that asking questions in the native language but instructing reasoning in English with “Let’s think step by step in English.” and “XLT” [65] indicates that translating questions into English and solving them step-by-step. “Trans” indicates that using the translation systems to convert questions into English and then solving them step-by-step. The experimental results demonstrate that the overall performance of the CoT instruction in English (En-CoT) is better than the one with instruction in native languages (Native-CoT). The results demonstrate that the CoT is still effective while using English to form the instructions for non-English queries. On the other hand, multilingual CoT attempts to enhance the reasoning capabilities of LLMs across multiple languages [66]. The multilingual CoT approach is especially beneficial for complex reasoning tasks deeply embedded in specific cultural contexts, enabling more natural and intuitive problem-solving [67, 68]. The common practice of multilingual CoT is to prompt the LLMs to establish a step-by-step reasoning process in the original language of queries, which can preserve linguistic and cultural nuances. In addition, the results corresponding to different sizes of LLaMA-2-Chat attend to examine the impact of scaling law with different multilingual inference strategies.
@@
The understanding of the syntax and semantics of various languages, such as idiomatic expressions and culture-specific references is challenging for existing LLMs. A recent study [69] introduces a cross-lingual CoT reasoning instruction fine-tuning framework, which randomly replaces language fragments with those from low-resource languages, and mixes the original and target languages within a single query. Besides, this study creates multilingual CoT instruction training data, which can be used to supervise fine-tuning LLMs to reduce the performance gap across different languages. However, developing datasets for training and evaluating multilingual CoT capabilities requires a robust representation of linguistic diversity and detailed language-specific nuances, which should be further addressed.
@@ 4.4 Code-Switching
Module Model SA [335] LID [335] MT [336] Summarization [337]
@@
Code-switching refers to the phenomenon where communicators switch between two or more languages during linguistic interactions, based on contextual needs. This phenomenon is common in bilingual or multilingual communities, especially in spoken communication [338, 339]. For example, in conversations involving Chinese-English or Spanish-English direction, speakers may freely switch between languages depending on the fluency, precision of expression, or emotional needs of the dialogue [340, 341, 335]. Solving code-switching texts is an important and challenging task as the language IDs are not specified before LLMs inference [70, 71].
@@
As shown in Table 3, we investigate the performance of LLMs on code-switching tasks. The results demonstrate that open-source LLMs cannot deal with the code-switching problem that needs to process multiple languages simultaneously under a single query. Multiple languages increase the complexity of processing and understanding language for models and LLMs need to continue to be supervised fine-tuned to accommodate simultaneous processing in this scenario. In addition, to improve the ability to deal with code-switching texts, researchers develop new pre-training techniques to enhance the understanding capacity of this linguistic phenomenon in LLMs [72]. For instance, Das et al. [342] modify the MLM [271] to enhance model reasoning performance in code-switching environments. By employing code-switching corpora, it masks tokens at the boundaries between two languages, thus forcing the model to learn semantic nuances at the points of language transition. Additionally, this work introduces modifications to the model structure by adding residual connections from intermediate layers to the final layer. An auxiliary loss function based on the representations of intermediate layers is also proposed, which enhances the cross-lingual ability to understand multiple languages in a single sentence. However, the cost of annotated code-switching data is expensive and it is challenging to investigate the non-parametric or semi-parametric method to adapt the code-switching problems.
4.5 Multilingual Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) is a methodology that integrates text generation with external knowledge retrieval, dynamically enhancing the quality of model response and accuracy by accessing relevant information [343, 344]. This approach enables the model to utilize up-to-date or specialized knowledge in text generation, thereby increasing its practicality and reliability. A main branch of the multilingual RAG adopts to retrieve knowledge from the open domain and applies it to the in-context (i.e., augmented prompt) [73, 74, 75, 76, 77, 78, 79, 80, 81]. In particular, Thakur et al. [82] propose NoMIRACL, a dataset across 18 languages, to evaluate the hallucination of LLM when given a piece of text in external retrieved knowledge. It consists of two parts, the non-relevant and the relevant subset, corresponding to the question and the passage which are non-related and related, respectively. The model is expected to answer “I don’t know” in the non-relevant subset, otherwise it is determined as a hallucination. In addition, hallucinations and off-target issues occur when incorporating LLMs with low-resource machine translation [83], where the RAG can mitigate these issues via improving the translation quality for low-resource directions [84, 85, 86, 87]. Overall, the RAG can effectively facilitate LLMs to generate more reliable responses, alleviating the issues of hallucination and factual error without fine-tuning. However, achieving substantial enhancements solely through the RAG method for low-resource languages, where LLMs struggle, poses a considerable challenge. Meanwhile, it is also challenging to build retrievers that can be used in low-resource languages [88].